# Accelerating Network Slice Placement with Multi-Agent Reinforcement Learning

### v1

@@ V-B Placement Algorithms
1) ILP Algorithm. Solves the optimization problem in (6) to obtain optimal placement but is computationally expensive. We use it as an upper bound on placement performance. A simplified version is employed by OSM MANO [38].
@@
5) Load Balancing Algorithm (Kubernetes Scheduler). The default Kubernetes scheduler, which places VNFs based on resource availability to balance utilization across the multi-cloud infrastructure [39].
@@ VI RELATED WORK
The problem of joint resource allocation and slice placement has been extensively studied and is often modeled as a variant of the virtual network embedding problem [9]. Several surveys address slice lifecycle management and placement optimization [1, 40], covering techniques for resource allocation, performance guarantees, and network efficiency. Optimization-based methods provide optimal decisions under latency, isolation, and resource constraints [11, 12], but are computationally prohibitive for large-scale deployments. Heuristic approaches reduce complexity [13, 15] but struggle with dynamic network conditions and lack real-world validation. More recently, reinforcement learning has been used to improve placement scalability and adaptability [18, 19, 20, 16], though most studies remain limited to simulation settings, hindering their applicability in real 5G systems.
Despite 5G-compliant testbeds supporting large-scale slice deployment [41, 21, 42], an optimization engine for intelligent slice placement remains absent. Existing efforts focus primarily on scaling slices rather than optimizing their placement [43, 44]. For example, [43] introduces a slice-as-a-service framework and evaluates deployments on local testbeds with cloud cost estimations, but without optimizing placement. Similarly, [44] explores slice scaling across various topologies and use cases, yet lacks mechanisms for intelligent placement. While these studies validate slice deployment and scaling, they overlook optimization under real-world constraints. To our knowledge, no existing framework jointly addresses slice placement, traffic profiling, resource demand estimation, and RL-based optimization in a real heterogeneous multi-cloud 5G setup—highlighting a key gap in the literature.
@@ VII CONCLUSION
This paper presents SlicePilot, a comprehensive framework for intelligent slice placement in heterogeneous multi-cloud infrastructures. Built on three functional blocks, SlicePilot seamlessly integrates (i) the User Plane Traffic Capturing Block to profile and replicate real-world 5G traffic patterns, (ii) the User Plane VNF Stressing Block to accurately estimate resource demands across slices, and (iii) the Slice Placement Block, a disaggregated multi-agent RL scheduler that automates VNF deployment while minimizing costs and ensuring SLA compliance. Through extensive experimentation on a multi-cloud testbed, SlicePilot delivers near-optimal placement efficiency, achieving deployment costs within 7.8% of ILP, 19× faster execution, and comparable SLA compliance, with only 2.42× more violations under high-load conditions.

### v2

@@ V-B Placement Algorithms
1) ILP Algorithm. Solves the optimization problem in (6) to obtain optimal placement but is computationally expensive. We use it as an upper bound on placement performance. A simplified version is employed by OSM MANO [osm].
@@
5) Load Balancing Algorithm (Kubernetes Scheduler). The default Kubernetes scheduler, which places VNFs based on resource availability to balance utilization across the multi-cloud infrastructure [k8s_scheduler].
@@ VI RELATED WORK
The problem of joint resource allocation and slice placement has been extensively studied and is often modeled as a variant of the virtual network embedding problem [surv6]. Several surveys address slice lifecycle management and placement optimization [surv1, surv3], covering techniques for resource allocation, performance guarantees, and network efficiency. Optimization-based methods provide optimal decisions under latency, isolation, and resource constraints [opt1, opt2], but are computationally prohibitive for large-scale deployments. Heuristic approaches reduce complexity [opt3, opt4] but struggle with dynamic network conditions and lack real-world validation. More recently, reinforcement learning has been used to improve placement scalability and adaptability [rl1, rl2, rl3, rl4], though most studies remain limited to simulation settings, hindering their applicability in real 5G systems.
Despite 5G-compliant testbeds supporting large-scale slice deployment [testbed1, testbed2, testbed3], an optimization engine for intelligent slice placement remains absent. Existing efforts focus primarily on scaling slices rather than optimizing their placement [scal2, scal3]. For example, [scal2] introduces a slice-as-a-service framework and evaluates deployments on local testbeds with cloud cost estimations, but without optimizing placement. Similarly, [scal3] explores slice scaling across various topologies and use cases, yet lacks mechanisms for intelligent placement. While these studies validate slice deployment and scaling, they overlook optimization under real-world constraints. To our knowledge, no existing framework jointly addresses slice placement, traffic profiling, resource demand estimation, and RL-based optimization in a real heterogeneous multi-cloud 5G setup—highlighting a key gap in the literature.
@@ VII CONCLUSION
This paper presents SlicePilot, a comprehensive framework for intelligent slice placement in heterogeneous multi-cloud infrastructures. Built on three functional blocks, SlicePilot seamlessly integrates (i) the User Plane Traffic Capturing Block to profile and replicate real-world 5G traffic patterns, (ii) the User Plane VNF Stressing Block to accurately estimate resource demands across slices, and (iii) the Slice Placement Block, a disaggregated multi-agent RL scheduler that automates VNF deployment while minimizing costs and ensuring SLA compliance. Through extensive experimentation on a self-managed multi-cloud testbed, SlicePilot delivers near-optimal placement efficiency, achieving deployment costs within 7.8% of ILP, 19× faster execution, and comparable SLA compliance, with only 2.42× more violations under high-load conditions. Unlike heuristic approaches, SlicePilot intelligently balances workload distribution, preventing costly edge overutilization while avoiding central cloud congestion, ensuring efficient, scalable, cost-effective deployments.